# Nonlinear Distortion Radiated from Large Arrays and Active Reconfigurable Intelligent Surfaces

### v1

@@ Abstract
Extremely large aperture arrays (ELAAs) and reconfigurable intelligent surfaces (RISs) are candidate enablers to realize connectivity goals for the sixth-generation (6G) wireless networks. For instance, ELAAs can provide orders-of-magnitude higher area throughput compared to what massive multiple-input multiple-output (MIMO) can deliver through spatial multiplexing, while RISs can improve the propagation conditions over wireless channels but a passively reflecting RIS must be large to be effective. Active RIS with amplifiers can deal with this issue. In this paper, we study the distortion created by nonlinear amplifiers in both ELAAs and active RIS. We analytically obtain the angular directions and depth of the nonlinear distortion in both near- and far-field channels. The results are demonstrated numerically and we conclude that non-linearities can both create in-band and out-of-band distortion that is beamformed in entirely new directions and distances from the transmitter.
@@ I Introduction
On the other hand, RISs can turn wireless propagation channels into programmable smart environments by software-controlled reflection of impinging signals [3, 4]. The aim can be to boost the received signal power at the receiver and improve other desired channel properties such as the MIMO channel rank. The basic idea is to densify the next generation of wireless networks with RISs instead of base stations (BSs) to create a low-cost, low-power, and high-throughput communication infrastructure. This new architecture is characterized by being programmable and reconfigurable, not only at the infrastructure nodes and user devices, but also in the surrounding wireless environment, properly shaped by programmable RISs. Each RIS acts as an advanced kind of transparent full-duplex relay/repeater [5]. Each antenna element of a RIS can be made of metamaterial, acts as an isotropic scatterer when it is sub-wavelength-sized, and the impedance can be tuned by diodes to create an inhomogeneous phase-shift pattern over the surface that reflects an incident wave as a beam in a desirable direction [6].
A conventional passive RIS reflects signals without amplification. This has the benefit of low power consumption but requires a large aperture size since the pathloss is the product of the pathlosses to and from the array, while the array gain grows quadratically with the aperture area [6]. Nevertheless, in practice, the capacity gains are typically only substantial in setups where the direct link between transmitter and receiver is completely blocked or very weak [7]. To expand the usefulness of the RIS technology to scenarios where the direct link is not weak, in [8], a new RIS architecture called active RISs was proposed to overcome the large path loss of reflected links. The key feature of active RISs is their ability to actively reflect signals with amplification. This can be realized by integrating a reflection-type amplifier into each reflecting element. Different from conventional relays, an active RIS remains to be transparent and full duplex.
@@ I-A Relevant Prior Art
The most relevant multi-antenna works are [14, 12, 15]. When it comes to MIMO arrays, only uniform linear arrays (ULAs) have been analyzed operating in the far-field region. Specifically, [14] showed that in single-user beamforming, nonlinear distortion is beamformed into the same direction as the desired signal, while in multi-user beamforming, the distortion is beamformed into many distinct directions different from those of the desired signals. Moreover, [12] showed that, if the input signal to the amplifiers has a dominant beam, the distortion is beamformed in the same way as that beam. When a superposition of multiple beams is transmitted without anyone being dominant, the distortion is practically isotropic. The transmission over frequency-selective channels gives rise to similar phenomena [15].
@@ I-B Contributions
Although the behavioral modeling of distortion has been studied in the multi-antenna literature, to the best of our knowledge, its effects on ELAAs and active RISs operating in the near-field region have not been addressed before, except in the preliminary version [1] of this manuscript, which studied only active RISs but operating in the far-field region. In this paper, we analyze the distortion either transmitted by an ELAA or created by an active RIS that reflects multiple signals. In both cases, the arrays are configured as uniform planar arrays (UPAs) and we consider operation in both the radiative near-field and far-field regions. Our main contributions can be summarized as follows:
@@
  • We derive a theorem that analytically characterizes the angular directions as well as the depth of nonlinear distortion in the near-field region. The theory can be used in system design to predict the effects of the nonlinear distortion and to take distortion effects into account when scheduling users in space and frequency.
  • We refine the number of nonlinear distortion directions that were derived in [12] for the far-field. Specifically, it was stated in [12, Th. 2] that third-degree non-linear distortion is beamformed in $(K^{3}-K^{2}+2K)/2$ directions. However, we show that this number is an upper bound and the exact number of directions can be substantially lower than that. Our results hold for both the near-field and far-field regions.
@@ II System Model
For later use, we introduce the $M\times 1$ vector notation of the signals at the RIS before and after amplification at time sample $n$ as ${\mathbf{x}}_{n}=\left[{x}_{1}[n],\ldots,{x}_{M}[n]\right]^{T}$ and ${\mathbf{y}}_{n}=\left[{y}_{1}[n],\ldots,{y}_{M}[n]\right]^{T}$, respectively. For analytical tractability, we consider a third-order memoryless polynomial distortion model [^2] Notice, that a third-order model is a reasonably good model for many hardware components in practice [14]. for the PAs at the ELAA and RIS. Specifically,
$\mathcal{A}\left({x}_{m}[n]\right)=\beta_{1}{x}_{m}[n]+\beta_{3}{x}_{m}[n]\left|{x}_{m}[n]\right|^{2},$ (2)
where $\beta_{1},\beta_{3}\in\mathbb{C}$ are the parameters for the first- and third-order term at the antenna elements.
@@ II-A Near- and Far-field Channel Models
We will now consider the radiative near-field and utilize the Fresnel region of the array response vector $\mathbf{a}(\varphi,\theta,r)$ between a UPA of $N$ identical antennas and a point of interest $\bm{\psi}=(\varphi,\theta,r)$. Here, $\varphi$ and $\theta$ can specify either the incoming wave arrival direction from a user located at the point $\bm{\psi}$ or the outgoing wave direction from the UPA towards the point $\bm{\psi}$ in terms of azimuth and elevation angles, respectively. Additionally, the value $r$ is the distance between the first element of the UPA (i.e., $m=0$) and the point of interest $\bm{\psi}$. The array response vector $\mathbf{a}(\bm{\psi})\in\mathbb{C}^{M\times 1}$ can be expressed as
@@ II-B Linear Precoded Signal
At the ELAA, the data symbols for the $K$ users are mapped to the antenna array by a precoder. We denote the transmitted symbol on subcarrier $l$ and time sample $n$ to user $k\ (k=1,\ldots,K)$ as $s_{k}[n,l]$, while the corresponding $K\times 1$ data symbols vector is denoted as $\mathbf{s}[n,l]=[s_{1}[n,l],\ldots,s_{K}[n,l]]^{T}$.
We assume that the ELAA has access to perfect channel state information (CSI), i.e., it has perfect knowledge of the realizations of the frequency-domain channel matrices $\{{\mathbf{H}}_{l}\}$ for $l\in\mathcal{S}$. We make this assumption since the goal is to characterize the impact of nonlinear distortion. If the CSI is imperfect, the signals will be transmitted toward the wrong locations and the distortion will be determined by these believed user locations rather than the exact ones, but our results otherwise hold.
Notice that, in general, the channel at a given subcarrier is frequency flat and therefore the data symbols are precoded individually for each subcarrier by the precoder $\mathbf{P}_{l}=\mathcal{P}({\mathbf{H}}_{l})$ which is frequency flat and a function $\mathcal{P}:\mathbb{C}^{K\times M}\rightarrow\mathbb{C}^{M\times K}$ of the channel. That is, the $M\times 1$ precoded symbol at time sample $n$ is given by $\mathbf{P}_{l}\mathbf{s}[n,l]$. However, since the channel is line-of-sight (LoS), its frequency response is constant over frequency, $\mathbf{H}=\mathbf{H}_{l}$ for $l\in\mathcal{S}$, and therefore the precoder becomes identical for all subcarriers, i.e., $\mathbf{P}_{l}=\mathbf{P}$ for $l\in\mathcal{S}$. The channel matrix is given by
$\mathbf{H}=\left[\mathbf{a}\left(\varphi_{1},\theta_{1},r_{1}\right),\dots,\mathbf{a}\left(\varphi_{K},\theta_{K},r_{K}\right)\right]^{T},$ (8)
where $\mathbf{a}(\cdot)$ is the array response vector defined in (3), while the ELAA observes the $k$-th user from the azimuth angle $\varphi_{k}$, elevation angle $\theta_{k}$ and distance $r_{k}$.
Then, the OFDM time-domain vector $\mathbf{x}_{n}$ before amplification at time sample $n$ is obtained from the data symbols as
$\mathbf{x}_{n}={\mathbf{P}}\mathbf{s}_{n},$ (9)
where
$\mathbf{s}_{n}\triangleq\frac{1}{\sqrt{N}}\sum_{l\in\mathcal{S}}\mathbf{s}[n,l]e^{jl\frac{2\pi}{N}n},$
for $n=0,\dots,N-1$ represents the time-domain data symbols vector. The transmitted signal before the amplification $\{\mathbf{x}_{n}\}$ satisfies the average transmit power constraint
@@
Let $\alpha_{mk}e^{j\phi_{m}\left(\varphi_{k},\theta_{k},r_{k}\right)}$ be the $(m,k)$-th entry of $\mathbf{P}$, i.e., the value of the precoder at the $m$-th antenna intended for the $k$-th user, where $\alpha_{mk}\in\mathbb{R}$ and $\phi_{m}(\cdot,\cdot,\cdot)\in\mathbb{R}$ denote the amplitude and phase of the precoder at the $(m,k)$-th entry. Then, from (9), the transmitted symbol before amplification at the $m$-th antenna element can be expressed as
$x_{m}[n]=\sum_{k=1}^{K}\alpha_{mk}e^{-j\frac{2\pi}{\lambda}\phi_{m}\left(\varphi_{k},\theta_{k},r_{k}\right)}s_{k}[n],$ (11)
where $s_{k}[n]$ is the $k$-th entry of the vector $\mathbf{s}_{n}$, i.e., the time-domain symbol intended for the $k$-th user. For simplicity, we assume the maximum-ratio transmition (MRT) precoder. On this basis, since we assume a LoS channel, the phase of the precoder $\phi_{m}(\cdot,\cdot,\cdot)$ is a conjugate match of the channel’s phase. In this case, the transmitted signal $x_{m}[n]$ is focused at the points $(\varphi_{k},\theta_{k},r_{k})$ for $k=1,\dots,K$.
@@ III Nonlinear Distortion Analysis
We are interested in the spatial characteristics of the third-order nonlinear term (or distortion) in (2) radiated from an ELAA. In the next theorem, we analytically derive the Fresnel approximation for the azimuth angle, elevation angle, and range of the near-field third-order distortion.
@@
Suppose the baseband signal in (11) is passed through the non-linear function in (2). Then, the radiative near-field radiation pattern of the third-order non-linear distortion is beamformed in a multitude of directions $(\varphi_{pqv},\theta_{pqv},r_{pqv})$ approximated analytically by
$\begin{split}\theta_{pqv}\approx\arcsin\left(\sin\theta_{p}-\sin\theta_{q}+\sin\theta_{v}\right)\end{split}$ (12)
[…]
$\begin{split}r_{pqv}\approx\left(\frac{1}{r_{p}}-\frac{1}{r_{q}}+\frac{1}{r_{v}}\right)^{-1}\end{split}$ (14)
for $p,q,v=1,\ldots,K$, where $\varphi_{pqv}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, $\theta_{pqv}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $r_{pqv}\geq 0$ are the azimuth, elevation and range of each focal point, respectively.
@@
The error of the Fresnel approximation used in Theorem 1 is small when $-2\tilde{\phi}_{m}(\varphi,\theta,r)/r\leq 0.1745$, which results in errors below $3.5\cdot 10^{-3}$[19],[20, cf. expression (12)]. In practice, this can be achieved when the distance between the ELAA and a user is approximately at least 6 meters. Theorem 1 can also be generalized when $K_{\tt FF}$ $(0<K_{\tt FF}\leq K)$ of the served users are in the far-field. With the next corollary we consider the extreme case where all users are in the far-field region.
@@
Moreover, notice that the focal points $(\varphi_{pqv},\theta_{pqv},r_{pqv})$ given in Theorem 1 are not unique. For instance, it easy to see that $(\varphi_{ppq},\theta_{ppq},r_{ppq})=(\varphi_{vvq},\theta_{vvq},r_{vvq})$, $\forall p,q,v$. Thus, with the next corollary, we provide all the unique focal points derived in Theorem 1.
@@
It is worth mentioning that when the argument of the function $\arcsin(\cdot)$ in (12) and (13) belongs to the subset of real numbers: $\left(-\infty,-1\right)\cup\left(1,\infty\right)$, then its corresponding value $(\theta_{pqv}$ or $\varphi_{pqv}$) is a complex number. In this case, the direction lacks a physical meaning and thus, it can be ignored since there is no physical location where the distortion can arise. Similarly, we consider only non-negative values of the range $r_{pqv}$. Therefore, the number of distortion directions of each subset in Corollary 2 are at most $|\mathcal{P}_{1}|\leq K$, $|\mathcal{P}_{2}|\leq K(K-1)$ and $|\mathcal{P}_{3}|\leq K(K-1)(K-2)/2$ which implies that the total number of unique focal points are at most $|\mathcal{P}|\leq(K^{3}-K^{2}+2K)/2$. This is in contrast to [12, Th. 2], where it is stated that the third-degree distortion is beamformed exactly into $(K^{3}-K^{2}+2K)/2$ directions.
@@
With the next corollaries, we highlight two special cases of how the distortion is radiated in the azimuth and elevation domain in the $yz$-plane.
@@
Suppose all users are located in the same elevation direction (i.e., $\theta_{k}=\theta_{o}$, $\forall k$) but different azimuth angles. The distortion is then focused at the points $(\varphi_{pqv},\theta_{pqv},r_{pqv})$:
@@
Suppose all users are located in the same azimuth direction (i.e., $\varphi_{k}=\varphi_{o}$, $\forall k$) but different elevation angles. Then, the distortion is beamformed into the directions $(\varphi_{pqv},\theta_{pqv},r_{pqv})$:
@@
Finally, it is worth mentioning that in the special case where the elevation angle $\theta_{o}$ in Corollary 3 and the azimuth angle $\varphi_{o}$ in Corollary 4 are zero, then the distortion’s directions in (19) and (20), respectively, become identical with those of a ULA and therefore these corollaries coincide with [12, Th. 2].

### v2

@@ Abstract
Extremely large aperture arrays (ELAAs) and reconfigurable intelligent surfaces (RISs) are candidate enablers to realize connectivity goals for the sixth-generation (6G) wireless networks. For instance, ELAAs can provide orders-of-magnitude higher area throughput compared to what massive multiple-input multiple-output (MIMO) can deliver through spatial multiplexing, while RISs can improve the propagation conditions over wireless channels but a passively reflecting RIS must be large to be effective. Active RIS with amplifiers can deal with this issue. In this paper, we analyze the distortion generated by nonlinear amplifiers in both ELAAs and active RIS. We derive analytical expressions for the angular directions and depth of nonlinear distortion in both near-field and far-field channels. These insights are then used in a distortion-aware scheduling scheme that predicts the beamforming directions of the distortion and strategically allocates users in frequency to minimize its impact. Numerical results validate our theoretical analysis and compare distortion-aware and distortion-unaware scheduling methods, highlighting the benefits of accounting for nonlinearities. We conclude that nonlinearities can both create in-band and out-of-band distortion that is beamformed in entirely new directions and distances from the transmitter.
@@ I Introduction
On the other hand, RISs can turn wireless propagation channels into programmable smart environments by software-controlled reflection of impinging signals [3, 4]. The aim can be to boost the received signal power at the receiver and improve other desired channel properties such as the MIMO channel rank. The basic idea is to densify the next generation of wireless networks with RISs instead of base stations (BSs) to create a low-cost, low-power, and high-throughput communication infrastructure. This new architecture is characterized by being programmable and reconfigurable, not only at the infrastructure nodes and user devices, but also in the surrounding wireless environment, properly shaped by programmable RISs. Each RIS acts as an advanced kind of transparent full-duplex relay/repeater [5, 6, 7]. Each unit cell of a RIS can be made of metamaterial, acts as an isotropic scatterer when it is sub-wavelength-sized, and the impedance can be tuned by diodes to create an inhomogeneous phase-shift pattern over the surface that reflects an incident wave as a beam in a desirable direction [8].
A conventional passive RIS reflects signals without amplification. This has the benefit of low power consumption but requires a large aperture size since the pathloss is the product of the pathlosses to and from the array, while the array gain grows quadratically with the aperture area [8]. Nevertheless, in practice, the capacity gains are typically only substantial in setups where the direct link between transmitter and receiver is completely blocked or very weak [9]. To expand the usefulness of the RIS technology to scenarios where the direct link is not weak, in [10], a new RIS architecture called active RISs was proposed to overcome the large path loss of reflected links. The key feature of active RISs is their ability to actively reflect signals with amplification. This can be realized by integrating a reflection-type amplifier into each reflecting cell. Different from conventional relays, an active RIS remains to be transparent and full duplex.
@@ I-A Relevant Prior Art
The most relevant multi-antenna works are [16, 13, 17]. When it comes to MIMO arrays, only uniform linear arrays (ULAs) have been analyzed operating in the far-field region. Specifically, [16] showed that in single-user beamforming, nonlinear distortion is beamformed into the same direction as the desired signal, while in multi-user beamforming, the distortion is beamformed into many distinct directions different from those of the desired signals. Moreover, [13] showed that, if the input signal to the amplifiers has a dominant beam, the distortion is beamformed in the same way as that beam. When a superposition of multiple beams is transmitted without anyone being dominant, the distortion is practically isotropic. The transmission over frequency-selective channels gives rise to similar phenomena [17]. Other studies [18, 19] highlight how nonlinear distortion critically affects MIMO systems and the trade-offs it imposes between system capacity and power efficiency. These insights reinforce the need to develop theoretical models that analyze the joint spatial and frequency behavior of such impairments, as done in this work, so the performance trade-offs can be rigorously optimized.
@@ I-B Contributions
Although the behavioral modeling of distortion has been studied in the multi-antenna literature, to the best of our knowledge, its effects on ELAAs and active RISs operating in the near-field region have not been addressed before, except in the preliminary version [1] of this manuscript, which studied only active RISs but operating in the far-field region. In this paper, we theoretically analyze the distortion introduced by an ELAA during transmission or by an active RIS during signal reflection. For both cases, the arrays are modeled as uniform planar arrays (UPAs), and their operation is examined in both the radiative near-field and far-field regions. Our main contributions can be summarized as follows:
@@
  • We derive a theorem that analytically characterizes the angular directions and depth of nonlinear distortion in the near-field region. This theoretical framework provides a foundation for developing signal processing strategies to mitigate distortion, thereby improving the reliability and performance of the communication system. To illustrate its practical utility, we present a scenario where the theorem is used to predict distortion beamforming directions and adjust user scheduling in frequency and space to reduce its impact.
  • We refine the number of nonlinear distortion directions that were derived in [13] for the far-field. Specifically, it was stated in [13, Th. 2] that third-degree non-linear distortion is beamformed in $(K^{3}-K^{2}+2K)/2$ directions. However, we show that this number is an upper bound and the exact number of directions can be substantially lower than that. Our results hold for both the near-field and far-field regions.
@@ II System Model
For later use, we introduce the $M\times 1$ vector notation of the signals before and after amplification at time sample $n$ as ${\mathbf{x}}_{n}=\left[{x}_{1}[n],\ldots,{x}_{M}[n]\right]^{T}$ and ${\mathbf{y}}_{n}=\left[{y}_{1}[n],\ldots,{y}_{M}[n]\right]^{T}$, respectively. The discrete baseband transfer function of the nonlinear PA is modeled by a memory polynomial and its output is related to its input by
$\mathcal{A}\left(x_{m}[n]\right)=\sum_{p=0}^{P}\sum_{l=0}^{L}\beta_{2p+1}[l]x_{m}[n-l]\left|x_{m}[n-l]\right|^{2p},$ (2)
where $\{\beta_{2p+1}\}\in\mathbb{C}$ are the polynomial parameters for the PA.
@@
[…]
Figure 1: The 3D geometry of an RIS consisting of $M_{y}$ elements per row and $M_{z}$ elements per column.
@@ II-A Near- and Far-field Channel Models
We will now consider the radiative near-field and utilize the Fresnel region of the array response vector $\mathbf{a}(\varphi,\theta,r)$ between a UPA of $M$ identical antennas and a point of interest $\bm{\psi}=(\varphi,\theta,r)$. Here, $\varphi$ and $\theta$ can specify either the incoming wave arrival direction from a user located at the point $\bm{\psi}$ or the outgoing wave direction from the UPA towards the point $\bm{\psi}$ in terms of azimuth and elevation angles, respectively. Additionally, the value $r$ is the distance between the first element of the UPA (i.e., $m=0$) and the point of interest $\bm{\psi}$. The array response vector $\mathbf{a}(\bm{\psi})\in\mathbb{C}^{M\times 1}$ can be expressed as
@@
Furthermore, the line-of-sight spherical wave channel model in the near-field region between the base station and user $k$ at the $\nu$-th subcarrier can be presented as[23, 24, 25]
^$\hat{}\mathbf{h}_{k}[\nu]=g_{k}\mathbf{a}(\varphi_{k},\theta_{k},r_{k})e^{-j2\pi\tau_{k}\nu},$ (8)
where $g_{k}$ and $\tau_{k}$ are the complex path gain and delay of the signal from the first antenna at the base station to user $k$, respectively. On this basis, the MIMO channel matrix $\hat{\mathbf{H}}_{\nu}\in\mathbb{C}^{M\times K}$ at the $\nu$-th subcarrier can be expressed as
$\hat{\mathbf{H}}_{\nu}=\underbrace{\left[{\mathbf{a}}_{1}\ {\mathbf{a}}_{2}\ \dots\ {\mathbf{a}}_{K}\right]}_{\triangleq\mathbf{A}}\mathbf{F}_{\nu},$ (9)
where ${\mathbf{a}}_{k}\triangleq\mathbf{a}(\varphi_{k},\theta_{k},r_{k})$ is the channel (steering) vector between the base station and $k$-th user and the $K\times K$ matrix $\mathbf{F}_{\nu}$ is diagonal with its $k$-th diagonal entry being $g_{k}e^{-j2\pi\tau_{k}\nu}$.
@@ II-B Linear Precoded Signal
At the ELAA, the data symbols for the $K$ users are mapped to the antenna array by a precoder. We denote the transmitted symbol on subcarrier $\nu$ to user $k\ (k=1,\ldots,K)$ as $s_{k}[\nu]$, while the corresponding $K\times 1$ data symbols vector is denoted as $\mathbf{s}_{\nu}=[s_{1}[\nu],\ldots,s_{K}[\nu]]^{T}$.
We assume that the ELAA has access to perfect channel state information (CSI), i.e., it has perfect knowledge of the realizations of the frequency-domain channel matrices {^$\{{\hat{}\mathbf{H}}_{\nu}\}$ for $\nu\in\mathcal{S}$. We make this assumption since the goal is to characterize the impact of nonlinear distortion. If the CSI is imperfect, the signals will be transmitted toward the wrong locations and the distortion will be determined by these believed user locations rather than the exact ones, but our results otherwise hold.
In general, the channel at a given subcarrier is frequency flat. Consequently, the data symbols are individually precoded for each subcarrier using the precoder $\mathbf{P}_{\nu}=\mathcal{P}({\hat{}\mathbf{H}}_{\nu})$, which is frequency flat and defined as a function $\mathcal{P}:\mathbb{C}^{K\times M}\rightarrow\mathbb{C}^{M\times K}$ of the channel. Specifically, the $M\times 1$ frequency-domain precoded symbol on subcarrier $\nu$ is expressed as ^$\hat{}\mathbf{z}_{\nu}=\mathbf{P}_{\nu}\mathbf{s}_{\nu}$. The OFDM time-domain vector $\mathbf{x}_{n}$ at time sample $n$, before amplification, is obtained by applying the IDFT to the precoded frequency-domain symbols and is given by:
$\mathbf{x}_{n}=\frac{1}{\sqrt{N}}\sum_{\nu\in\mathcal{S}}{\mathbf{P}_{\nu}}\mathbf{s}_{\nu}e^{j\nu\frac{2\pi}{N}n},$ (10)
for $n=0,...,N-1$. In essence, the frequency-domain precoded symbol ^$\hat{}\mathbf{z}_{\nu}$ is transformed into the time domain via an IDFT.
Common linear precoders are the maximum-ratio transmission (MRT) and the zero-forcing (ZF), given by the following expressions:
$\mathcal{P}_{\tt MRT}\left(\hat{\mathbf{H}}_{\nu}\right)$ $=\alpha{\mathbf{A}^{\ast}}\mathbf{F}_{\nu}^{\ast},$ (11)
$\mathcal{P}_{\tt ZF}\left(\hat{\mathbf{H}}_{\nu}\right)$ $=\alpha\mathbf{A}^{\ast}\underbrace{\left(\mathbf{F}_{\nu}\mathbf{A}^{T}\mathbf{A}^{\ast}\right)^{-1}}_{\triangleq\mathbf{B}_{\nu}},$ (12)
respectively, where the matrix $\mathbf{B}_{\nu}\in\mathbb{C}^{K\times K}$ is defined for later use. Furthermore, $\alpha$ is a constant used for power normalization that is chosen such that the transmitted signal before the amplification $\{\mathbf{x}_{n}\}$ satisfies the average transmit power constraint
@@
Note that in the expressions (11) and (12), we assume a multi-user (MU) scenario where all $K$ co-scheduled users occupy the same set of subcarriers. With MRT, the resulting time-domain vector $\mathbf{x}_{n}$ simplifies to:
$\mathbf{x}_{n}={\mathbf{A}^{\ast}}\mathbf{s}_{n},$ (14)
where $\mathbf{s}_{n}\triangleq({1}/{\sqrt{N}})\sum_{\nu\in\mathcal{S}}\alpha\mathbf{F}_{\nu}^{\ast}\mathbf{s}_{\nu}\exp\left\{j2\pi\nu n/N\right\}$. It is important to note that when using the ZF, we obtain a similar expression to (14), except that the vector $\mathbf{s}_{n}$ depends on $\mathbf{B}_{\nu}$.
Let $e^{j\phi_{m}\left(\varphi_{k},\theta_{k},r_{k}\right)}$ represent the $(m,k)$-th entry of $\mathbf{A}^{\ast}$, where $\phi_{m}\left(\varphi_{k},\theta_{k},r_{k}\right)\in\mathbb{R}$ is the phase of the precoder, respectively. Then, from (14), the transmitted symbol before amplification at the $m$-th antenna element can be expressed as
$x_{m}[n]=\sum_{k=1}^{K}e^{-j\frac{2\pi}{\lambda}\phi_{m}\left(\varphi_{k},\theta_{k},r_{k}\right)}s_{k}[n],$ (15)
where $s_{k}[n]$ is the $k$-th entry of the vector $\mathbf{s}_{n}$. Notice that since we assume a line-of-sight (LoS) channel, the phase of the precoder $\phi_{m}(\cdot,\cdot,\cdot)$ is a conjugate match of the channel’s phase. In this case, the transmitted signal $x_{m}[n]$ is focused at the points $(\varphi_{k},\theta_{k},r_{k})$ for $k=1,\dots,K$.
[…]
@@ III Nonlinear Distortion Analysis
We are interested in the spatial characteristics of the nonlinear term (or distortion) in (2) radiated from an ELAA. In the following theorem, we derive the Fresnel approximation for the azimuth angle, elevation angle, and range of the $p$-th order distortion in the near-field.
@@
Suppose the baseband signal in (15) is passed through the non-linear function defined in (2). In that case, the radiative near-field radiation pattern of the $p$-order non-linear distortion is beamformed in multiple directions $(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})$ which can be approximated analytically as follows:
$\theta_{\bm{k}}$ $\approx\arcsin\left(\sum_{i=0}^{2p}(-1)^{i}\sin\theta_{k_{i}}\right),$ (16)
$\varphi_{\bm{k}}$ $\approx\arcsin\left(\frac{1}{\cos\theta_{\bm{k}}}\sum_{i=0}^{2p}(-1)^{i}\sin\varphi_{k_{i}}\cos\theta_{k_{i}}\right),$ (17)
$r_{\bm{k}}$ $\approx\left(\sum_{i=0}^{2p}(-1)^{i}\frac{1}{r_{k_{i}}}\right)^{-1}.$ (18)
where ${\bm{k}}\triangleq\left(k_{0},k_{1},\dots,k_{2p}\right)$ with ${\bm{k}}\in\{1,2,\ldots,K\}^{2p}$, $\varphi_{\bm{k}}\in[-\frac{\pi}{2},\frac{\pi}{2}]$, $\theta_{\bm{k}}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and $r_{\bm{k}}\geq 0$ are the azimuth, elevation and range of each focal point, respectively.
@@
The error of the Fresnel approximation used in Theorem 1 is small when $-2\tilde{\phi}_{m}(\varphi,\theta,r)/r\leq 0.1745$, which results in errors below $3.5\cdot 10^{-3}$[26],[27, cf. expression (12)]. In practice, this can be achieved when the distance between the ELAA and a user is approximately at least 6 meters.
Remark 1.
The derivation of Theorem 1 assumes that all users allocate the same set of subcarriers. However, the results of the theorem also hold in the case of sub-band scheduling, where different users allocate distinct portions of the bandwidth. In this scenario, the transmit signal for each user can still be expressed in the form shown in (14), with appropriate modifications to $\mathbf{s}_{n}$ to account for the allocation of different subcarriers. This ensures that the key principles and conclusions derived in the theorem remain valid under sub-band scheduling.
Theorem 1 can also be generalized when $K_{\tt FF}$ $(0<K_{\tt FF}\leq K)$ of the served users are in the far-field. With the next corollary we consider the extreme case where all users are in the far-field region.
@@
Moreover, notice that the focal points $(\varphi_{\bm{k}},\theta_{\bm{k}},r_{\bm{k}})$ given in Theorem 1 are not unique. For instance, it is easy to see that for the third-order non-linear distortion $(\varphi_{ppq},\theta_{ppq},r_{ppq})=(\varphi_{vvq},\theta_{vvq},r_{vvq})$, $\forall p,q,v$. Thus, with the next corollary, we provide all the unique focal points derived in Theorem 1.
@@
Notice that the $p$-th focal point in $\mathcal{P}_{1}$ represents the distortion beamformed at the location of the $p$-th user, where, as expected, $|\mathcal{P}_{1}|=K$. It can be observed from (49) that these focal points carry more power than those in $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$, which implies that distortion is expected to be stronger in the direction of the users. Furthermore, it is worth mentioning that when the argument of the function $\arcsin(\cdot)$ in (16) and (17) belongs to the subset of real numbers: $\left(-\infty,-1\right)\cup\left(1,\infty\right)$, then its corresponding value $(\theta_{pqv}$ or $\varphi_{pqv}$) is a complex number. In this case, the direction lacks a physical meaning and thus, it can be ignored since there is no physical location where the distortion can arise. Similarly, we consider only non-negative values of the range $r_{pqv}$. Therefore, the number of distortion directions of $\mathcal{P}_{2}$ and $\mathcal{P}_{3}$ in Corollary 2 are at most $|\mathcal{P}_{2}|\leq K(K-1)$ and $|\mathcal{P}_{3}|\leq K(K-1)(K-2)/2$ which implies that the total number of unique focal points are at most $|\mathcal{P}|\leq(K^{3}-K^{2}+2K)/2$. This is in contrast to [13, Th. 2], where it is stated that the third-degree distortion is beamformed exactly into $(K^{3}-K^{2}+2K)/2$ directions.
@@
With the next corollaries, we highlight two special cases of how the third-order distortion is radiated in the azimuth and elevation domain in the $yz$-plane.
@@
Suppose all users are located in the same elevation direction (i.e., $\theta_{k}=\theta_{o}$, $\forall k$) but different azimuth angles. The third-order distortion is then focused at the points $(\varphi_{pqv},\theta_{pqv},r_{pqv})$:
@@
Suppose all users are located in the same azimuth direction (i.e., $\varphi_{k}=\varphi_{o}$, $\forall k$) but different elevation angles. Then, the third-order distortion is beamformed into the directions $(\varphi_{pqv},\theta_{pqv},r_{pqv})$:
@@
Finally, it is worth mentioning that in the special case where the elevation angle $\theta_{o}$ in Corollary 3 and the azimuth angle $\varphi_{o}$ in Corollary 4 are zero, then the distortion’s directions in (23) and (24), respectively, become identical with those of a ULA and therefore these corollaries coincide with [13, Th. 2].